# A Quality Diversity Method to Automatically Generate Multi-Agent Path Finding Benchmark Maps

### v3

@@ 1 Introduction
We study the problem of generating diverse and targeted benchmark maps for Multi-Agent Path Finding (MAPF) algorithms. Given a map and a group of agents, MAPF is the problem of finding collision-free paths from their start to goal locations. MAPF has wide applications in coordinating hundreds of robots in automated warehouses (Li et al. 2021b; Varambally, Li, and Koenig 2022; Zhang et al. 2023a, b), moving characters in video games (Ma et al. 2017; Jansen and Sturtevant 2008), managing drone traffic (Ho et al. 2022; Choudhury et al. 2022), and controlling multi-robotic arms (Shaoul et al. 2024).
@@
Given the wide applicability of MAPF, many algorithms are proposed to solve MAPF, and benchmarking these algorithms becomes an important task. Stern et al. (2019) have proposed a set of 33 MAPF benchmark maps, covering a diverse spectrum of map sizes, layouts, and difficulties. [^1] https://movingai.com/benchmarks/mapf/index.html These human-designed maps are widely used to test newly proposed MAPF algorithms (Li, Ruml, and Koenig 2021; Li et al. 2021a; Okumura et al. 2019; Skrynnik et al. 2024). However, such benchmark maps have several problems. First, they may not fully encompass the failure modes of an algorithm. Second, with a small fixed set of maps, it is challenging to sufficiently understand the pros and cons of different algorithms in different maps. Third, while comparing multiple algorithms, researchers are prone to run experiments on a subset of maps that are biased towards the proposed algorithms, resulting in an unfair comparison.
@@
Meanwhile, Quality Diversity (QD) algorithms have been used to generate a diverse set of high-quality solutions by optimizing a given objective function and a set of diversity measure functions. A recent work (Zhang et al. 2023b) has used QD algorithms to optimize layouts for automated warehouses. To ensure that the optimized layouts possess human-explainable and regularized patterns, a follow-up work (Zhang et al. 2023a) then leverages Neural Cellular Automata (NCA) to generate the layouts, while applying QD algorithms to optimize the parameters of the NCA. Cellular Automata (Gardner 1970) iteratively generates complex tile-based structures from a simple one through local interaction between tiles. Each tile decides its next state based on its neighbors via a fixed rule. Neural Cellular Automata (NCA) represents the rule using a convolutional neural network.
@@
In this paper, we adapt the layout optimization approach based on the QD algorithm and NCA from the previous work (Zhang et al. 2023a) and use it with an alternative goal of generating diverse benchmark maps for MAPF algorithms. To demonstrate that our approach can generate benchmark maps for a broad spectrum of MAPF algorithms, we choose five algorithms, namely CBS (Sharon et al. 2015), EECBS (Li, Ruml, and Koenig 2021), PBS (Ma et al. 2019), PIBT (Okumura et al. 2019), and Learn-to-Follow (LTF) (Skrynnik et al. 2024), representing search-based, priority-based, rule-based, and learning-based algorithms, respectively.
@@ 2.1 Multi-Agent Path Finding (MAPF)
Search-Based. This category includes exponential-time algorithms that exhaustively explore the solution space of MAPF. They usually have theoretical guarantees such as optimality and suboptimality, but suffer from computational time. Examples include M (Wagner and Choset 2011, 2015), BCP (Lam et al. 2019), ICTS (Sharon et al. 2013), CBS (Sharon et al. 2015; Gange, Harabor, and Stuckey 2019; Li et al. 2020), ECBS (Barer et al. 2014) and EECBS (Li, Ruml, and Koenig 2021). We choose CBS and EECBS as the representatives of this category. CBS starts by planning a shortest path for each agent, ignoring the collisions, and then resolving the collisions with a two-level search. The high-level search iteratively selects unresolved collisions and adds constraints to tackle them. The low-level search runs single-agent planning to compute new paths that satisfy the constraints. The search continues until all collisions are resolved. EECBS is a variant of CBS that uses Explicit Estimation Search (Thayer and Ruml 2011) on the high level and focal search (Pearl and Kim 1982) on the low level.
@@
Priority-Based. Priority-based algorithms plan paths for each agent in a sequence, forcing agents with low priority to avoid colliding with those with high priority. The priority planning (PP) (Erdmann and Lozano-Pérez 1987) algorithm plans paths for agents in a pre-defined priority order. Monte-Carlo PP (Friedrich et al. 2024) pre-defines a set of randomized priority orders and selects the paths with the best solutions. PBS (Ma et al. 2019) combines CBS with PP to explicitly search for a good priority order with a two-level search. The high-level search of PBS is similar to CBS except that PBS constrains one agent to have a higher priority order than the other. The low-level search then plans paths with the order from the high-level search. We choose PBS as the representative of this category.
@@
Rule-Based. Rule-based algorithms (Wang and Botea 2011; Okumura 2023) leverage pre-defined rules to move agents to their goals. They usually run much faster but produce worse solutions than search-based and priority-based methods. We choose PIBT (Okumura et al. 2019) as the representative of this category. PIBT uses an iterative one-timestep rule to move the agents. At each timestep, each agent plans a single-step action towards its goal. In case of collision, a pre-defined rule based on priority inheritance and backtracking is applied to resolve the collision.
@@
Learning-Based. Learning-based algorithms usually formulate MAPF as a Multi-Agent Reinforcement Learning (MARL) (Sartoretti et al. 2019; Damani et al. 2021) problem. They train a shared policy for each agent, taking its local field of view as input and deciding its next action. We choose Learn to Follow (LTF) (Skrynnik et al. 2024) as the representative of this category. LTF starts by searching for a guide path for each agent without considering collisions. Then it uses a shared learnable policy to move the agents to their goals along the guide paths while avoiding collisions. LTF is developed for lifelong MAPF, a variant of MAPF that constantly assigns new goals to agents. We modify it for MAPF by asking agents to stop when they reach their first goals.
@@ 2.2 MAPF Benchmarks
Early works in MAPF typically use randomly generated benchmark maps (Wagner and Choset 2015) or maps from the single-agent path planning benchmark set (Sturtevant 2012). Stern et al. (2019) established the major set of MAPF benchmark maps with 33 maps of five categories (game, maze, empty, random, and room) and map sizes ranging from 32 $\times$ 32 to 256 $\times$ 256. This benchmark set is used extensively in MAPF research (Okumura et al. 2019; Li et al. 2021a; Okumura 2023; Friedrich et al. 2024). However, while testing novel MAPF algorithms, instead of using all maps, researchers usually select a subset of them of different categories and sizes. This could result in unfair and biased benchmarking results because they can intentionally or unintentionally cherry-pick maps that favor their proposed algorithms, or the design of their algorithms could overfit to the selected maps.
The MARL community usually generates random maps to test the trained policy for MAPF problems (Sartoretti et al. 2019; Damani et al. 2021). Researchers who develop MAPF algorithms for the multi-robot systems for automated warehouses or sortation centers often use a set of warehouse maps (Li et al. 2021b; Varambally, Li, and Koenig 2022; Zhang et al. 2023b, a, 2024).
@@ 2.3 Quality Diversity (QD) Algorithms
QD algorithms (Cully et al. 2015; Lehman and Stanley 2011a, b) are inspired by evolutionary algorithms with diversity search to generate a diverse collection of high-quality solutions by optimizing an objective function and diversifying a set of diversity measure functions simultaneously. QD algorithms maintain an archive, which is a tessellated measure space defined by the measure functions. The archive stores the best solution in each tessellated cell. QD algorithms then optimize the sum of objective values of all solutions in the archive, defined as QD-score. We choose Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine and Nikolaidis 2023) as the method for generating benchmark maps because it is the state-of-the-art QD algorithm specialized for continuous search domains. CMA-MAE is an extension of MAP-Elites (Mouret and Clune 2015) that incorporates the covariance matrix adaptation mechanism of CMA-ES (Hansen 2016), which is a derivative-free single-objective optimizer. CMA-ES maintains a multi-variate Gaussian distribution and iteratively samples from it for new solutions. It evaluates the solutions and updates the Gaussian toward high-objective regions. CMA-MAE adapts this mechanism to optimize the QD-score.
@@ 2.4 Automatic Scenario Generation
Automated methods of generating scenarios to understand, test, or benchmark a given algorithm have been proposed by different fields. In autonomous driving, researchers have generated scenarios to test and evaluate developed autonomous driving systems (Arnold and Alexander 2013; Abeysirigoonawardena, Shkurti, and Dudek 2019; Mullins et al. 2018). In human-robot-interaction, prior works have used QD algorithms to generate shared autonomy scenarios (Fontaine et al. 2021b) or diverse kitchen layouts (Fontaine et al. 2021a) to understand the coordination behavior between humans and robots. A follow-up work (Bhatt et al. 2023) leverages model-based QD methods (Bhatt et al. 2022; Zhang et al. 2022) to improve the sample efficiency of QD algorithms. In reinforcement learning, prior works have generated maps to benchmark (Cobbe et al. 2020) or continuously improve (Wang et al. 2019, 2020) trained agent.
In MAPF, one recent work (Ren et al. 2024) has used QD to generate maps for MAPF. Our work differs in two key aspects: First, they focus on generating maps of fixed difficulties, quantified by an approximate metric based on map connectivity, while we are interested in generating benchmarking maps targeted for specific MAPF algorithms. Second, they directly optimize tile types (i.e., whether it is an empty space or an obstacle) using MAP-Elites, while we optimize a map generator based on NCA, which has been shown to be more effective for generating maps with diverse regularized patterns (Zhang et al. 2023a).
@@ 3 Benchmark Generation Approach
We adapt the previous work (Zhang et al. 2023a) to use CMA-MAE and NCA to generate diverse benchmark maps with the objective and measures computed by running MAPF algorithms. Figure 1 shows the overview of the method. Starting by sampling $b$ parameter vectors $\boldsymbol{\theta}$, we form $b$ NCA generators and use them to generate $b$ maps. Maps generated by NCA might not be valid. We then adapt a mixed integer linear programming (MILP) solver (Zhang et al. 2020, 2023a, 2023b) to repair the map such that (1) the map is valid, and (2) the number of obstacles falls in a pre-defined range $[O_{lb},O_{ub}]$. We then run MAPF algorithms to evaluate the objective and measure values and add the evaluated maps to an archive. For each evaluation, we run the given MAPF algorithm in $N_{e}$ instances. Finally, we update the Gaussian distribution, starting a new iteration. We repeat this process until we evaluate $N_{eval}$ maps.
MAPF Instance. In this paper, we focus on generating MAPF benchmark maps of a pre-defined size and number of agents. To generate a MAPF instance based on a map, we follow the bucket method (Stern et al. 2019) to generate evenly distributed start and goal locations of the agents with a distance constraint, which was used for generating the “even” scenarios in the MAPF benchmark. For each map, we generate $N_{e}$ MAPF instances with different starts and goals, run MAPF algorithms on each instance, and compute the average objective and measures of all instances as the objective and measures of the map.
@@ 3.1 Objective
Two-Algorithm Experiments. To compare two given MAPF algorithms in two-algorithm experiments, we aim to generate maps that maximize the performance gap between the two algorithms. We consider two pairs of comparisons: (1) EECBS vs. PBS, and (2) PIBT vs. LTF. We compare EECBS and PBS because (1) both suboptimal algorithms use a two-level search based on collision avoidance, (2) both find near-optimal solutions empirically, and (3) few prior works have compared them thoroughly. We compare PIBT and LTF because (1) both use a pre-defined policy, either rule-based or learned, to move the agents step by step to their goals while avoiding collisions, (2) both run fast but suffer from deadlocks, and (3) while the LTF work (Skrynnik et al. 2024) has compared them on warehouse maps, showing that LTF outperforms PIBT, we want to find cases where PIBT outperforms LTF.
@@ 3.2 Diversity Measures
General Hardness. To measure the general hardness of a map, the most commonly used metric is the number of obstacles. The number of obstacles in a map affects the space available to resolve collisions as well as the search space of the algorithms. Therefore, diversifying it would give a range of difficulties. We also consider two metrics that are proposed specifically to reflect the hardness of a map: the standard deviation of Betweenness Centrality (std of BC) (Ewing et al. 2022) and $\lambda_{2}$ (Ren et al. 2024). To compute the std of BC, we search for the shortest path between every pair of empty spaces in the map and then calculate the standard deviation of the usage of all empty spaces. The higher the std of BC, the more congestion and the harder the MAPF problem. To compute $\lambda_{2}$ of a map, we view the map as a graph $G(V,E)$, and compute the second smallest eigenvalue of the normalized Laplacian matrix of $G$. Ren et al. (2024) shows empirically that $\lambda_{2}$ is correlated with the hardness of the map.
Spatial Arrangement. Aside from the hardness of the maps, we want to diversify the spatial arrangement of the maps, making the generated maps stylistically diverse. We first consider the KL divergence of the tile pattern distribution (Fontaine et al. 2021b) between the tile pattern distribution of the generated map and one category of fixed maps selected from the MAPF benchmark (Stern et al. 2019). A tile pattern is one possible arrangement of the obstacles and empty spaces in a 3 $\times$ 3 grid. To compute the tile pattern distribution of a map, we count the number of all possible tile patterns in the map. We then compute the KL divergence between the tile pattern distribution of the map and a fixed distribution computed from all the maze maps [^2] https://movingai.com/benchmarks/maze/index.html of the human-designed MAPF benchmark set (Stern et al. 2019). The smaller the KL divergence of tile pattern distribution, the more similar in local patterns our generated map is to the selected set of benchmark maps. We choose maze maps because they empirically yield the most complex patterns. In addition, we consider the entropy of the tile pattern distribution (Zhang et al. 2023a). The lower the entropy, the more regularized patterns the map possesses. Finally, we consider the KL divergence of the Weisfeiler-Lehman (WL) graph feature (Shervashidze et al. 2011) between the generated map and the maze maps. Similar to the tile pattern distribution, a smaller KL divergence of the WL graph feature implies more similarities between the local patterns of the generated map and the maze maps.

### v4

@@ 1 Introduction
We study the problem of generating diverse and targeted benchmark maps for Multi-Agent Path Finding (MAPF) algorithms. Given a map and a group of agents, MAPF is the problem of finding collision-free paths from their start to goal locations. MAPF has wide applications in coordinating hundreds of robots in automated warehouses Li et al. (2021c); Zhang et al. (2023a, b), moving characters in video games Ma et al. (2017), managing drone traffic Choudhury et al. (2022), and controlling multi-robotic arms Shaoul et al. (2024).
@@
Given the wide applicability of MAPF, many algorithms are proposed to solve MAPF, and benchmarking these algorithms becomes an important task. Stern et al. (2019) have proposed a set of 33 MAPF benchmark maps, covering a diverse spectrum of map sizes, layouts, and difficulties [^1] https://movingai.com/benchmarks/mapf/index.html. These human-designed maps are widely used to test newly proposed MAPF algorithms Li et al. (2021b, a); Okumura et al. (2019); Skrynnik et al. (2024). However, such benchmark maps have several problems. First, they may not fully encompass the failure modes of an algorithm. Second, with a small fixed set of maps, it is challenging to sufficiently understand the pros and cons of different algorithms in different maps. Third, while comparing multiple algorithms, researchers are prone to run experiments on a subset of maps that are biased towards the proposed algorithms, resulting in an unfair comparison.
@@
Meanwhile, Quality Diversity (QD) algorithms have been used to generate a diverse set of high-quality solutions by optimizing a given objective function and a set of diversity measure functions. A recent work Zhang et al. (2023b) has used QD algorithms to optimize layouts for automated warehouses. To ensure that the optimized layouts possess human-explainable and regularized patterns, a follow-up work Zhang et al. (2023a) then leverages Neural Cellular Automata (NCA) to generate the layouts, while applying QD algorithms to optimize the parameters of the NCA. Cellular Automata Gardner (1970) iteratively generates complex tile-based structures from a simple one through local interaction between tiles. Each tile decides its next state based on its neighbors via a fixed rule. Neural Cellular Automata (NCA) represents the rule using a convolutional neural network.
@@
In this paper, we adapt the layout optimization approach based on the QD algorithm and NCA from the previous work Zhang et al. (2023a) and use it with an alternative goal of generating diverse benchmark maps for MAPF algorithms. To demonstrate that our approach can generate benchmark maps for a broad spectrum of MAPF algorithms, we choose five algorithms, namely CBS Sharon et al. (2015), EECBS Li et al. (2021b), PBS Ma et al. (2019), PIBT Okumura et al. (2019), and Learn-to-Follow (LTF) Skrynnik et al. (2024), representing search-based, priority-based, rule-based, and learning-based algorithms, respectively.
@@ 2.1 Multi-Agent Path Finding (MAPF)
Search-Based. This category includes exponential-time algorithms that exhaustively explore the solution space of MAPF. They usually have theoretical guarantees such as optimality and suboptimality, but suffer from computational time. Examples include M Wagner and Choset (2015), BCP Lam et al. (2019), CBS Sharon et al. (2015); Li et al. (2020), ECBS Barer et al. (2014) and EECBS Li et al. (2021b). We choose CBS and EECBS as the representatives of this category. CBS starts by planning a shortest path for each agent, ignoring the collisions, and then resolving the collisions with a two-level search. The high-level search iteratively selects unresolved collisions and adds constraints to tackle them. The low-level search runs single-agent planning to compute new paths that satisfy the constraints. The search continues until all collisions are resolved. EECBS is a suboptimal variant of CBS that uses Explicit Estimation Search Thayer and Ruml (2011) on the high level and focal search Pearl and Kim (1982) on the low level.
@@
Priority-Based. Priority-based algorithms plan paths for each agent in a sequence, forcing agents with low priority to avoid colliding with those with high priority. The priority planning (PP) Erdmann and Lozano-Pérez (1987) algorithm plans paths for agents in a pre-defined priority order. Monte-Carlo PP Friedrich et al. (2024) pre-defines a set of randomized priority orders and selects the paths with the best solutions. PBS Ma et al. (2019) combines CBS with PP to explicitly search for a good priority order with a two-level search. The high-level search of PBS is similar to CBS except that PBS constrains one agent to have a higher priority order than the other. The low-level search then plans paths with the order from the high-level search. We choose PBS as the representative of this category.
@@
Rule-Based. Rule-based algorithms leverage pre-defined rules to move agents to their goals. They usually run much faster but produce worse solutions than search-based and priority-based methods. We choose PIBT Okumura et al. (2019) as the representative of this category. PIBT uses an iterative one-timestep rule to move the agents. At each timestep, each agent plans a single-step action towards its goal. In case of collision, a pre-defined rule based on priority inheritance and backtracking is applied to resolve the collision.
@@
Learning-Based. Learning-based algorithms usually formulate MAPF as a Multi-Agent Reinforcement Learning (MARL) Sartoretti et al. (2019) problem. They train a shared policy for each agent, taking its local field of view as input and deciding its next action. We choose Learn to Follow (LTF) Skrynnik et al. (2024) as the representative of this category. LTF starts by searching for a guide path for each agent without considering collisions. Then it uses a shared learnable policy to move the agents to their goals along the guide paths while avoiding collisions. LTF is developed for lifelong MAPF, a variant of MAPF that constantly assigns new goals to agents. We modify it for MAPF by asking agents to stop when they reach their first goals.
@@ 2.2 MAPF Benchmarks
Early works in MAPF typically use randomly generated benchmark maps Wagner and Choset (2015) or maps from the single-agent path planning benchmark set Sturtevant (2012). Stern et al. (2019) established the major set of MAPF benchmark maps with 33 maps, which is used extensively in MAPF research Li et al. (2021a); Okumura (2023); Friedrich et al. (2024). However, researchers tend to select a subset of all the maps while testing novel MAPF algorithms. This could result in unfair and biased benchmarking results because they can intentionally or unintentionally cherry-pick maps that favor their proposed algorithms, or the design of their algorithms could overfit to the selected maps. The MARL community usually generates random maps to test the trained policy for MAPF problems Sartoretti et al. (2019). Researchers who develop MAPF algorithms for the multi-robot systems for automated warehouses or sortation centers often use a set of warehouse maps Li et al. (2021c); Zhang et al. (2023b, a, 2024).
@@ 2.3 Quality Diversity (QD) Algorithms
QD algorithms Cully et al. (2015) are inspired by evolutionary algorithms with diversity search to generate a diverse collection of high-quality solutions by optimizing an objective function and diversifying a set of diversity measure functions. QD algorithms maintain an archive, which is a tessellated measure space defined by the measure functions. The archive stores the best solution in each tessellated cell. QD algorithms then optimize the sum of objective values of all solutions in the archive, defined as QD-score. We choose Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) Fontaine and Nikolaidis (2023) as the method for generating benchmark maps because it is the state-of-the-art QD algorithm specialized for continuous search domains. CMA-MAE is an extension of MAP-Elites Mouret and Clune (2015) that incorporates the covariance matrix adaptation mechanism of CMA-ES Hansen (2016), which is a derivative-free single-objective optimizer. CMA-ES maintains a multi-variate Gaussian distribution and iteratively samples from it for new solutions. It evaluates the solutions and updates the Gaussian toward high-objective regions. CMA-MAE adapts this mechanism to optimize the QD-score.
@@ 2.4 Automatic Scenario Generation
Automated methods of generating scenarios to understand, test, or benchmark a given algorithm have been proposed by different fields. In autonomous driving, researchers have generated scenarios to test and evaluate developed autonomous driving systems Abeysirigoonawardena et al. (2019). In human-robot-interaction, prior works have used QD algorithms to generate shared autonomy scenarios Fontaine et al. (2021b) or diverse kitchen layouts Fontaine et al. (2021a) to understand the coordination behavior between humans and robots. A follow-up work Bhatt et al. (2023) leverages model-based QD methods Bhatt et al. (2022); Zhang et al. (2022) to improve the sample efficiency of QD algorithms. In reinforcement learning, prior works have generated maps to benchmark Cobbe et al. (2020) or continuously improve Wang et al. (2019) trained agents.
In MAPF, one recent work Ren et al. (2024) has used QD to generate maps for MAPF. Our work differs in two key aspects: First, they focus on generating maps of fixed difficulties, quantified by an approximate metric based on map connectivity, while we are interested in generating benchmarking maps targeted for specific MAPF algorithms. Second, they directly optimize tile types (i.e., whether it is an empty space or an obstacle) using MAP-Elites, while we optimize a map generator based on NCA, which has been shown to be more effective for generating maps with diverse regularized patterns Zhang et al. (2023a).
@@ 3 Benchmark Generation Approach
We adapt the previous work Zhang et al. (2023a) to use CMA-MAE and NCA to generate diverse benchmark maps with the objective and measures computed by running MAPF algorithms. Figure 1 shows the overview of the method. Starting by sampling $b$ parameter vectors $\boldsymbol{\theta}$, we form $b$ NCA generators and use them to generate $b$ maps. Maps generated by NCA might not be valid. We then adapt a mixed integer linear programming (MILP) solver Zhang et al. (2023b) to repair the map such that (1) the map is valid, and (2) the number of obstacles falls in a pre-defined range $[O_{lb},O_{ub}]$. We then run MAPF algorithms to evaluate the objective and measure values and add the evaluated maps to an archive. For each evaluation, we run the given MAPF algorithm in $N_{e}$ instances. Finally, we update the Gaussian distribution, starting a new iteration. We repeat this process until we evaluate $N_{eval}$ maps.
MAPF Instance. In this paper, we focus on generating MAPF benchmark maps of a pre-defined size and number of agents. To generate a MAPF instance based on a map, we follow the bucket method Stern et al. (2019) to generate evenly distributed start and goal locations of the agents with a distance constraint, which was used for generating the even scenarios in the MAPF benchmark. For each map, we generate $N_{e}$ MAPF instances with different starts and goals, run MAPF algorithms on each instance, and compute the average objective and measures of all instances as the objective and measures of the map.
@@ 3.1 Objective
Two-Algorithm Experiments. To compare two given MAPF algorithms in two-algorithm experiments, we aim to generate maps that maximize the performance gap between the two algorithms. We consider two pairs of comparisons: (1) EECBS vs. PBS, and (2) PIBT vs. LTF. We compare EECBS and PBS because (1) both suboptimal algorithms use a two-level search based on collision avoidance, (2) both find near-optimal solutions empirically, and (3) few prior works have compared them thoroughly. We compare PIBT and LTF because (1) both use a pre-defined policy, either rule-based or learned, to move the agents step by step to their goals while avoiding collisions, (2) both run fast but suffer from deadlocks, and (3) while the LTF work Skrynnik et al. (2024) has compared them on warehouse maps, showing that LTF outperforms PIBT, we want to find cases where PIBT outperforms LTF.
@@ 3.2 Diversity Measures
General Hardness. To measure the general hardness of a map, the most commonly used metric is the number of obstacles. The number of obstacles in a map affects the space available to resolve collisions as well as the search space of the algorithms. Therefore, diversifying it would give a range of difficulties. We also consider two metrics that are proposed specifically to reflect the hardness of a map: the standard deviation of Betweenness Centrality (std of BC) Ewing et al. (2022) and $\lambda_{2}$ Ren et al. (2024). To compute the std of BC, we search for the shortest path between every pair of empty spaces in the map and then calculate the standard deviation of the usage of all empty spaces. The higher the std of BC, the more congestion and the harder the MAPF problem. To compute $\lambda_{2}$ of a map, we view the map as a graph $G(V,E)$, and compute the second-smallest eigenvalue of the normalized Laplacian matrix of $G$. Ren et al. (2024) shows empirically that $\lambda_{2}$ is correlated with the hardness of the map.
Spatial Arrangement. Aside from the hardness of the maps, we want to diversify the spatial arrangement of the maps, making the generated maps stylistically diverse. We first consider the KL divergence of the tile pattern distribution Fontaine et al. (2021b) between the tile pattern distribution of the generated map and one category of fixed maps selected from the MAPF benchmark Stern et al. (2019). A tile pattern is one possible arrangement of the obstacles and empty spaces in a 3 $\times$ 3 grid. To compute the tile pattern distribution of a map, we count the number of all possible tile patterns in the map. We then compute the KL divergence between the tile pattern distribution of the map and a fixed distribution computed from all the maze maps [^2] https://movingai.com/benchmarks/maze/index.html of the MAPF benchmark set Stern et al. (2019). The smaller the KL divergence of tile pattern distribution, the more similar in local patterns our generated map is to the selected set of benchmark maps. We choose maze maps because they empirically yield the most complex patterns. In addition, we consider the entropy of the tile pattern distribution Zhang et al. (2023a). The lower the entropy, the more regularized patterns the map possesses. Finally, we consider the KL divergence of the Weisfeiler-Lehman (WL) graph feature Shervashidze et al. (2011) between the generated map and the maze maps. Similar to the tile pattern distribution, a smaller KL divergence of the WL graph feature implies more similarities between the local patterns of the generated map and the maze maps.